OpenStudy (hhopke):

A food company distributes its tomato soup in two cans of different sizes. For the larger can, both the diameter and the height have been increased by 30% By what percentage does the volume of the can increase from the smaller can to the larger can? Round your answer to the nearest percent.

OpenStudy (anonymous):

Small can volume = (π*r^2) height Large can volume = (π*1.3r^2) height you can give r and height all the values you want, and compare them, to make it simple, we can try with 1and 2 SCV = (π*1^2) 2 = 6.2831 LCV = (π*1.3(1)^2) = 10.6185 Then you only have to calculate the percentage of the growth. 10.6185 - 6.2831 = 4.3354 6.2831= 100% 4.3354 = 68.96% Percentage of growth = 69 % ok, you could have guessed it was 69%, my result was 68.96 cause I didn't use all the decimal numbers I should have used.

OpenStudy (hhopke):

Okay. Thank you so much @Skielerlucas04!

OpenStudy (anonymous):

NP